A random vector (X1, X2) is a discrete random vector if its space D or range space is finite or countable. Suppose X1, X2 are positive integer valued and we need to find P(X1= x1, X2= x2), where x1, x2 are known positive integers, in terms of their known cumulative distribution function F. 

1
Expert's answer
2022-01-31T16:14:44-0500

"P(X_1= x_1, X_2= x_2)=P(X_1= x_1, X_2\leq x_2)-P(X_1= x_1, X_2< x_2)"

"=P(X_1= x_1, X_2\leq x_2)-P(X_1= x_1, X_2\leq x_2-1)"


"P(X_1= x_1, X_2\leq x_2)=P(X_1\leq x_1, X_2\leq x_2)-P(X_1< x_1, X_2\leq x_2)"

"=P(X_1\leq x_1, X_2\leq x_2)-P(X_1\leq x_1-1, X_2\leq x_2)"

"=F(x_1,x_2)-F(x_1-1,x_2)"


Similarly, "P(X_1= x_1, X_2\leq x_2-1)=F(x_1,x_2-1)-F(x_1-1,x_2-1)".

Therefore,

"P(X_1= x_1, X_2= x_2)="

"=(F(x_1,x_2)-F(x_1-1,x_2))-(F(x_1,x_2-1)-F(x_1-1,x_2-1))"

"=F(x_1,x_2)-F(x_1-1,x_2)-F(x_1,x_2-1)+F(x_1-1,x_2-1)"